# Prediction of two-neutron halos in the $N=28$ isotones ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$

### v1

@@ Abstract
The ground states of the nuclei ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ are investigated using the hyperspherical formalism. Since they are located at the edge of the “big island of inversion”, we concentrate on whether we are likely to find two-neutron Borromean halos in these nuclei. A three-body model with effective $n$-$n$ and ${{}^{38}\mathrm{Mg}}+n$ interactions is built for ${}^{40}$g based on the available data. We also give predictions for the low-lying spectrum of ${}^{38}\mathrm{Na}={{}^{37}\mathrm{Na}}+n$ and two-neutron separation energy of the ${}^{39}\mathrm{Na}$ nucleus. Depending on parameter choice, we report an increase in the matter radii in the range $0.1$ $0.5$ fm relative to those of the core nuclei. The results suggest a two-neutron halo structure in ${}^{40}\mathrm{Mg}$ for a subset of parameters, reinforcing the prediction of a Borromean halo nucleus. The calculations indicate that a two-neutron halo is even more likely for ${}^{39}\mathrm{Na}$. As expected, the halo is linked to the disappearance of the shell gap in these nuclei due to the inversion of the $2p_{3/2}$ and $1f_{7/2}$ orbitals. We study the total cross section for scattering of these nuclei from a carbon target using a Glauber model and show that these provide a clear signal to assess the halo structure.
preprint: NITEP 197
I Introduction
The development of novel radioactive ion beam facilities has made it possible to study the shell evolution across the nuclear chart, and, most interestingly, near the neutron drip line. Neutron-rich nuclides display exotic structural features such as dampening of shell gaps Warburton et al. (1990); Dobaczewski et al. (1994); Sorlin and Porquet (2008), shape coexistence Heyde and Wood (2011); Otsuka and Tsunoda (2016); Nishibata et al. (2019) and formation of halos and deformed structure Tanihata et al. (1985); Motobayashi et al. (1995); Nakamura et al. (2009); Takechi et al. (2012); Hamamoto (2012); Crawford et al. (2019); Bagchi et al. (2020).
@@ I Introduction
Above $N=20$, an explanation of the magic numbers requires a spin-orbit force Otsuka et al. (2020). Thus, $N=28$ is the smallest magic number that is sensitive to the spin-orbit force. As shown in the lower panel of Fig. 1, for $20\leq N\leq 28$ the $1f_{7/2}$ orbit is filled, while in the absence of spin-orbit splitting the $1f_{7/2}$ and $2p_{3/2}$ orbits are almost degenerate. With modifications to the spin-orbit force, the spacing between these two orbits may change, and this can lead to an “inverted” order, where the $2p_{3/2}$ lies below the $1f_{7/2}$. The same result may also be caused by deformation Warburton et al. (1990); Horiuchi et al. (2022). Such islands of inversion (IOI) have been found in isolated cases. It has also been shown that for $N=20$ and $28$ the IOIs merge in the neon, sodium, and magnesium isotopic chains, creating a “big island of inversion”, called the B-IOI Caurier et al. (2014); Doornenbal et al. (2013).
@@
The disappearance, or “melting”, of $N=28$ shell gap has been confirmed experimentally in neutron-rich Al, Si, S, and Ar isotopes Scheit et al. (1996); Glasmacher et al. (1997); Takeuchi et al. (2012); Sohler et al. (2002); Doornenbal et al. (2013); Nowak et al. (2016); Calinescu et al. (2016); Crawford et al. (2019); Momiyama et al. (2020); Longfellow et al. (2020). For $Z=11$, the lowest proton number in the B-IOI, the heaviest observed neutron-rich nucleus is ${}^{39}\mathrm{Na}$ Ahn et al. (2022); Zhang et al. (2023), but the possibility of the existence of even heavier Na isotopes has not yet been excluded. Very few studies Zhang et al. (2023); Suzuki and Kimura (2021); Suzuki et al. (2022) examine the melting of $N=28$ shell gap in Na and Mg isotones, among these only Zhang et al. (2023) delving into halo formation in ${}^{39}\mathrm{Na}$ using the deformed relativistic Hartree-Bogoliubov framework.
@@
The recent observation of the disappearance of the $N=20$ shell gap at the low-$Z$ side of the $N=20$ chain (shown in the red circle in the upper panel of Fig. 1), has led to the identification of the ${}^{29}\mathrm{F}$ system as the heaviest known two-neutron Borromean-halo nucleus Singh et al. (2020); Fortunato et al. (2020); Casal et al. (2020); Bagchi et al. (2020). Motivated by this observation, it is interesting to explore the low-$Z$ side of the $N=28$ shell closure for potential two-neutron Borromean halos in the Na and Mg isotopes (as shown in the blue circles in the upper panel of Fig. 1). The magnesium isotopic chain also provides interesting candidates to explore the transition from strongly-bound to weakly-bound nuclei and the form of $n$-$n$ correlations at the limits of stability. Deformation will also play a role: e.g., the nucleus ${}^{32}\mathrm{Mg}$, with $N=20$, can be described as a prolate-deformed rotor Motobayashi et al. (1995); Nowacki et al. (2021). There is only limited knowledge of the properties of heavier neutron-rich Mg isotopes with $N\geq 25$. The weakly-bound nucleus ${}^{37}\mathrm{Mg}$ ($N=25$) has been observed, and its ground state can be interpreted as a one-neutron halo with the valence neutron occupying $2p_{3/2}$ orbit Takechi et al. (2014); Kobayashi et al. (2014). The neighbouring odd-A nucleus ${}^{39}\mathrm{Mg}$ was confirmed as unbound, indicating the important role of pair correlations.
@@
For the present study, we start from the medium-mass open shell nucleus ${}^{40}\mathrm{Mg}$, which is the next two-neutron halo candidate after ${}^{29}\mathrm{F}$ according to various theoretical models  Hamamoto (2007); Caurier et al. (2014); Nakada and Takayama (2018); Macchiavelli et al. (2022). The evaluated (i.e., not measured but estimated based on systematics) two-neutron separation energy of ${}^{40}\mathrm{Mg}$ is $s_{2n}=0.670\pm 0.710\,\mathrm{MeV}$Wang et al. (2021). This nucleus was first observed experimentally in 2007 Baumann et al. (2007) and more recently low-lying excited states of ${}^{40}\mathrm{Mg}$ have been studied using a one-proton removal reaction from ${}^{41}\mathrm{Al}$ Crawford et al. (2019). Recent ab initio calculations suggest large deformation in ${}^{40}$g Yuan et al. (2024), though no connection to halo physics is made.
@@
Because ${}^{39}\mathrm{Mg}$ is neutron unbound, it is reasonable to assume that the correlation between the two valence neutrons in ${}^{40}\mathrm{Mg}$ plays an important role in binding this system. Thus, the ${}^{40}\mathrm{Mg}$ nucleus, when interpreted as a ${}^{38}\mathrm{Mg}$ core plus two valence neutrons, provides an example of a Borromean system, similar to the well-studied two-neutron halo nuclei ${}^{6}$e or ${}^{11}\mathrm{Li}$ Tanihata et al. (2013). Because three-body models give a good description of the structural properties of many Borromean nuclei, we apply the same technique, using effective three-body calculations to study the configuration mixing in and matter radius of the ground state of ${}^{40}\mathrm{Mg}$. Given the uncertainties in the two-neutron separation energy, we explore how the allowed values influence the configuration mixing and matter radius, by analysing different scenarios. This allows us to explore the possibility of halo formation in this region of the nuclear chart, well beyond the heaviest observed two-neutron halo in ${}^{29}\mathrm{F}$Singh et al. (2020); Fortunato et al. (2020); Casal et al. (2020); Bagchi et al. (2020). The most recent atomic mass evaluation Wang et al. (2021) indicates that ${}^{39}\mathrm{Na}$ is unbound, contradicting Ahn et al. recent experimental discovery; our calculations within same three-body framework aim to predict the two-neutron separation energy ($s_{2n}$) and matter radii for ${}^{39}\mathrm{Na}$ systematically. Finally, we also provide Glauber-model predictions for the total reaction cross sections of ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ on a carbon target, which will further support the finding of halo formation in these systems.
II Model Formulation
@@
To focus on the effective $\text{core}+n+n$ system, we employ the hyperspherical formalism Zhukov et al. (1993); Nielsen et al. (2001). The eigenstates of the three-body Hamiltonian can be written as
@@
Here, $\Upsilon_{Klm_{l}}^{l_{x}l_{y}}(\Omega)$ are the hyperspherical harmonics Zhukov et al. (1993), eigenstates of the hypermomentum operator $\widehat{K}$, and $\kappa_{S_{x}}$ is a spin function. This equation implies that $\bm{l}=\bm{l}_{x}+\bm{l}_{y}$ is the total orbital angular momentum, $S_{x}$ is the coupled spin of the two neutrons, and the total $j$ results from $\bm{j}=\bm{l}+\bm{S}_{x}$. Note that we assume the core to be inert and its spin to be zero. More details can be found, for instance, in Ref. Casal (2018).
@@
where $C_{i\beta}$ are the diagonalization coefficients to be determined numerically. For this purpose, different bases have been used in various works Descouvemont et al. (2003); Matsumoto et al. (2004); Rodríguez-Gallardo et al. (2005), and in the present work, we employ the analytical transformed harmonic oscillator (THO) basis from Ref. Casal et al. (2013).
The diagonalization of the three-body Hamiltonian requires the computation of the corresponding kinetic energy and potential matrix elements. With the above definition in hyperspherical coordinates, we can write Casal (2016); Thompson et al. (2004)
@@
In this expression, $V_{ij}$ are the two-body interactions within the three-body composite system, to be fixed from the known information on the binary subsystems, and $V_{\rm 3b}(\rho)$ is a phenomenological three-body force, to account for effects not included in the two-body interactions alone Thompson et al. (2004); Rodríguez-Gallardo et al. (2005); de Diego et al. (2010); Casal et al. (2013). This term contains the only free parameter in this model and is used to adjust the energy of the three-body states to the experimental values if known.
III Two-body ($\text{core}+n$) subsystems
@@
We will start with ${}^{38}\mathrm{Mg}+n$ system. Though $N=26$ has a partially filled $1f_{7/2}$ subshell, we model ${}^{39}\mathrm{Mg}$ as a ${}^{38}\mathrm{Mg}$ core surrounded by an unbound neutron moving in the $f_{7/2}$, $p_{3/2}$, $f_{5/2}$, and $p_{1/2}$ orbitals in a simple independent-particle shell-model picture. Even though the distinction between the core and valence neutrons is not completely obvious, for simplicity, we nevertheless make an inert-core approximation. This means that the effects from internal rearrangement or core-valence exchange must be contained in effective potential parameters. Note that a similar approach has been successfully followed in other three-body calculations, such as the ${}^{27}\text{F}+n$ and the ${}^{29}\text{F}+n$ subsystems for the description of ${}^{29}\mathrm{F}~{}$Singh et al. (2020); Fortunato et al. (2020); Casal et al. (2020) and ${}^{31}\mathrm{F}~{}$Singh et al. (2022). Core excitations could be incorporated by using, for instance, a particle-rotor model Macchiavelli et al. (2022) with an effective core deformation parameter.
The only theoretical study on the spectrum of the neutron-unbound nucleus ${}^{39}\mathrm{Mg}$ uses the Gamow shell model, and reports an indication of a low-lying resonance at $0.129$ MeV with $J^{\pi}$ either $7/2^{-}$ or $3/2^{-}$ Fossez et al. (2016). However, due to the unavailability of experimental data, this does not rule out the existence of other possibilities such as a single resonant structure or more than two states, and in any case, the spin-parity assignment is not unambiguous.
Table 1: Parameter sets for the ${}^{38}\mathrm{Mg}+n$ (upper-panel) and ${}^{37}\mathrm{Na}+n$ (lower-panel) Woods-Saxon interactions, Eq. (III). Here $a$ is diffuseness, $V^{(l)}_{0}$ is the potential depth and $E_{R}$ is the position of the resonances. Note that $r_{0}=1.25$ fm, and $V_{ls}=16.842$ MeV (for ${}^{38}\mathrm{Mg}+n$) and $16.324$ MeV (for ${}^{37}\mathrm{Na}+n$) are fixed.
@@
$V_{{\rm core}+n}(r)$ $=$ $\left(-V^{(l)}_{0}+V_{ls}\lambda_{\pi}^{2}\bm{l}\cdot\bm{s}\frac{1}{r}\frac{d}{dr}\right)$
×$\times\frac{1}{1+\exp[(r-R_{c})/a]}\,,$
where $V^{l}_{0}$ is in general $l$-dependent, and $R_{c}=r_{0}A_{c}^{1/3}$ with $A_{c}$ the mass number of the core. The spin-orbit interaction is written in terms of the Compton wavelength $\lambda_{\pi}=1.414$ fm. Following Ref. Horiuchi et al. (2010a), the spin-orbit strength is taken to follow the systematic trend Bohr and Mottelson (1969) and has the value $V_{ls}=16.842$ MeV for ${}^{38}\mathrm{Mg}+n$ and $16.324$ MeV for ${}^{37}\mathrm{Na}+n$. The value $r_{0}=1.25$ fm is as originally suggested for ${}^{31}\mathrm{Ne}$  Horiuchi et al. (2010a); Singh et al. (2020), and we examine three scenarios, see Table 1. In Set $1$ and $2$, $V_{0}$ is chosen to be $l$-independent. In set $1$ it is adjusted to fix the $f_{7/2}$ ground-state resonance of ${}^{39}\mathrm{Mg}$ at $0.129$ MeV, corresponding to the Gamow shell model energy Fossez et al. (2016). This set gives a “normal” shell-model scenario, with an additional $p_{3/2}$ resonance appearing at higher energy, $0.349$ MeV.
@@
These potentials generate Pauli forbidden single-particle states, which would result in unphysical eigenstates of the three-body Hamiltonian, and need to be removed. In this work, this is achieved through a supersymmetric transformation Baye (1987a, b).
@@
As we do not have either theoretical or experimental predictions and data for ${}^{38}\mathrm{Na}$, we use the same potential parameters as for ${}^{39}\mathrm{Mg}$. The only changes are $R_{c}$ and the spin-orbit strength, as explained above. The latter takes into account effectively the tensor force between the unpaired proton and valence neutron Hagino and Sagawa (2016); Otsuka et al. (2010). Indeed, since ${}^{39}\mathrm{Na}$ has an odd proton number, its ground state angular momentum is not zero. However, we have found before Singh et al. (2020); Casal and Garrido (2020); Singh et al. (2022) that approximating the ground state of these light systems as a $0^{+}$ state gives reasonable results. The lower half of Table 1 shows the parameter sets for the ${}^{37}\mathrm{Na}+n$ potential. Our predictions for the low-lying spectrum of ${}^{38}\mathrm{Na}$ show no “normal” ordering of the single-particle states; most scenarios present an inversion, indicating a shell-evolution effect, that we describe effectively by just changing the spin-orbit strength.
IV Three-body ($\text{core}+n+n$) calculations

### v2

@@ Abstract
The ground states of the nuclei ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ are investigated using the hyperspherical formalism. Since they are located at the edge of the “big island of inversion”, we concentrate on whether we are likely to find two-neutron Borromean halos in these nuclei. A three-body model with effective $n$-$n$ and ${{}^{38}\mathrm{Mg}}+n$ interactions is built for ${}^{40}\mathrm{Mg}$ based on the available data. We also give predictions for the low-lying spectrum of ${}^{38}\mathrm{Na}={{}^{37}\mathrm{Na}}+n$ and two-neutron separation energy of the ${}^{39}\mathrm{Na}$ nucleus. Depending on parameter choice, we report an increase in the matter radii in the range $0.1$ $0.5$ fm relative to those of the core nuclei. The results suggest a two-neutron halo structure in ${}^{40}\mathrm{Mg}$ for a subset of parameters, reinforcing the prediction of a Borromean halo nucleus. The calculations indicate that a two-neutron halo is even more likely for ${}^{39}\mathrm{Na}$. As expected, the halo is linked to the disappearance of the shell gap in these nuclei due to the inversion of the $2p_{3/2}$ and $1f_{7/2}$ orbitals. We study the total cross section for scattering of these nuclei from a carbon target using a Glauber model and show that these provide a clear signal to assess the halo structure.
journal: Physics Letters B
1 Introduction
The development of novel radioactive ion beam facilities has made it possible to study the shell evolution across the nuclear chart, and, most interestingly, near the neutron drip line. Neutron-rich nuclides display exotic structural features such as dampening of shell gaps [1, 2, 3], shape coexistence [4, 5, 6] and formation of halos and deformed structure [7, 8, 9, 10, 11, 12, 13].
@@ 1 Introduction
Above $N=20$, an explanation of the magic numbers requires a spin-orbit force [14]. Thus, $N=28$ is the smallest magic number that is sensitive to the spin-orbit force. As shown in the lower panel of Fig. 1, for $20\leq N\leq 28$ the $1f_{7/2}$ orbit is filled, while in the absence of spin-orbit splitting the $1f_{7/2}$ and $2p_{3/2}$ orbits are almost degenerate. With modifications to the spin-orbit force, the spacing between these two orbits may change, and this can lead to an “inverted” order, where the $2p_{3/2}$ lies below the $1f_{7/2}$. The same result may also be caused by deformation [1, 15]. Such islands of inversion (IOI) have been found in isolated cases. It has also been shown that for $N=20$ and $28$ the IOIs merge in the neon, sodium, and magnesium isotopic chains, creating a “big island of inversion”, called the B-IOI [16, 17].
@@
The disappearance, or “melting”, of $N=28$ shell gap has been confirmed experimentally in neutron-rich Al, Si, S, and Ar isotopes [18, 19, 20, 21, 17, 22, 23, 12, 24, 25]. For $Z=11$, the lowest proton number in the B-IOI, the heaviest observed neutron-rich nucleus is ${}^{39}\mathrm{Na}$ [26, 27], but the possibility of the existence of even heavier Na isotopes has not yet been excluded. Very few studies [27, 28, 29] examine the melting of $N=28$ shell gap in Na and Mg isotones, among these only [27] delving into halo formation in ${}^{39}\mathrm{Na}$ using the deformed relativistic Hartree-Bogoliubov framework.
@@
The recent observation of the disappearance of the $N=20$ shell gap at the low-$Z$ side of the $N=20$ chain (shown in the red circle in the upper panel of Fig. 1), has led to the identification of the ${}^{29}\mathrm{F}$ system as the heaviest known two-neutron Borromean-halo nucleus [30, 31, 32, 13]. Motivated by this observation, it is interesting to explore the low-$Z$ side of the $N=28$ shell closure for potential two-neutron Borromean halos in the Na and Mg isotopes (as shown in the blue circles in the upper panel of Fig. 1). The magnesium isotopic chain also provides interesting candidates to explore the transition from strongly-bound to weakly-bound nuclei and the form of $n$-$n$ correlations at the limits of stability. Deformation will also play a role: e.g., the nucleus ${}^{32}\mathrm{Mg}$, with $N=20$, can be described as a prolate-deformed rotor [8, 33]. There is only limited knowledge of the properties of heavier neutron-rich Mg isotopes with $N\geq 25$. The weakly-bound nucleus ${}^{37}\mathrm{Mg}$ ($N=25$) has been observed, and its ground state can be interpreted as a one-neutron halo with the valence neutron occupying $2p_{3/2}$ orbit [34, 35, 36, 37, 38]. The neighbouring odd-A nucleus ${}^{39}\mathrm{Mg}$ was confirmed as unbound, indicating the important role of pair correlations.
@@
For the present study, we start from the medium-mass open shell nucleus ${}^{40}\mathrm{Mg}$, which is the next two-neutron halo candidate after ${}^{29}\mathrm{F}$ according to various theoretical models  [39, 16, 40, 41]. The evaluated (i.e., not measured but estimated based on systematics) two-neutron separation energy of ${}^{40}\mathrm{Mg}$ is $s_{2n}=0.670\pm 0.710\,\mathrm{MeV}$[42]. This nucleus was first observed experimentally in 2007 [43] and more recently low-lying excited states of ${}^{40}\mathrm{Mg}$ have been studied using a one-proton removal reaction from ${}^{41}\mathrm{Al}$ [12]. Recent ab initio calculations suggest large deformation in 40Mg [44], though no connection to halo physics is made.
@@
Because ${}^{39}\mathrm{Mg}$ is neutron unbound, it is reasonable to assume that the correlation between the two valence neutrons in ${}^{40}\mathrm{Mg}$ plays an important role in binding this system. Thus, the ${}^{40}\mathrm{Mg}$ nucleus, when interpreted as a ${}^{38}\mathrm{Mg}$ core plus two valence neutrons, provides an example of a Borromean system, similar to the well-studied two-neutron halo nuclei 6He [45, 46] or ${}^{11}\mathrm{Li}$ [47]. Because three-body models give a good description of the structural properties of many Borromean nuclei, we apply the same technique, using effective three-body calculations to study the configuration mixing in and matter radius of the ground state of ${}^{40}\mathrm{Mg}$. Given the uncertainties in the two-neutron separation energy, we explore how the allowed values influence the configuration mixing and matter radius, by analysing different scenarios. This allows us to explore the possibility of halo formation in this region of the nuclear chart, well beyond the heaviest observed two-neutron halo in ${}^{29}\mathrm{F}$[30, 31, 32, 13]. The most recent atomic mass evaluation [42] indicates that ${}^{39}\mathrm{Na}$ is unbound, contradicting Ahn et al. [26] recent experimental discovery; our calculations within same three-body framework aim to predict the two-neutron separation energy ($s_{2n}$) and matter radii for ${}^{39}\mathrm{Na}$ systematically. Finally, we also provide Glauber-model predictions for the total reaction cross sections of ${}^{40}\mathrm{Mg}$ and ${}^{39}\mathrm{Na}$ on a carbon target, which will further support the finding of halo formation in these systems.
2 Model Formulation
@@
To focus on the effective $\text{core}+n+n$ system, we employ the hyperspherical formalism [48, 49]. The eigenstates of the three-body Hamiltonian can be written as
@@
Here, $\Upsilon_{Klm_{l}}^{l_{x}l_{y}}(\Omega)$ are the hyperspherical harmonics [48], eigenstates of the hypermomentum operator $\widehat{K}$, and $\kappa_{S_{x}}$ is a spin function. This equation implies that $\bm{l}=\bm{l}_{x}+\bm{l}_{y}$ is the total orbital angular momentum, $S_{x}$ is the coupled spin of the two neutrons, and the total $j$ results from $\bm{j}=\bm{l}+\bm{S}_{x}$. Note that we assume the core to be inert and its spin to be zero. More details can be found, for instance, in Ref. [50].
@@
where $C_{i\beta}$ are the diagonalization coefficients to be determined numerically. For this purpose, different bases have been used in various works [51, 52, 53], and in the present work, we employ the analytical transformed harmonic oscillator (THO) basis from Ref. [54]. The THO functions are built by performing a local scale transformation to the harmonic oscillator functions, changing their Gaussian asymptotic behaviour into an exponential decay, which is better suited to accurately describe bound nuclear states with a relatively small set of basis functions. This procedure was also adopted, for instance, in Refs. [30, 32].
The diagonalization of the three-body Hamiltonian requires the computation of the corresponding kinetic energy and potential matrix elements. With the above definition in hyperspherical coordinates, we can write [55, 56]
@@
In this expression, $V_{ij}$ are the two-body interactions within the three-body composite system, to be fixed from the known information on the binary subsystems, and $V_{\rm 3b}(\rho)$ is a phenomenological three-body force, to account for effects not included in the two-body interactions alone [56, 53, 57, 54]. This term contains the only free parameter in this model and is used to adjust the energy of the three-body states to the experimental values if known.
3 Two-body ($\text{core}+n$) subsystems
@@
We will start with ${}^{38}\mathrm{Mg}+n$ system. Though $N=26$ has a partially filled $1f_{7/2}$ subshell, we model ${}^{39}\mathrm{Mg}$ as a ${}^{38}\mathrm{Mg}$ core surrounded by an unbound neutron moving in the $f_{7/2}$, $p_{3/2}$, $f_{5/2}$, and $p_{1/2}$ orbitals in a simple independent-particle shell-model picture. Even though the distinction between the core and valence neutrons is not completely obvious, for simplicity, we nevertheless make an inert-core approximation. This means that the effects from internal rearrangement or core-valence exchange must be contained in effective potential parameters. Note that a similar approach has been successfully followed in other three-body calculations, such as the ${}^{27}\text{F}+n$ and the ${}^{29}\text{F}+n$ subsystems for the description of ${}^{29}\mathrm{F}~{}$[30, 31, 32] and ${}^{31}\mathrm{F}~{}$[58]. Core excitations could be incorporated by using, for instance, a particle-rotor model [41] with an effective core deformation parameter.
The only theoretical study on the spectrum of the neutron-unbound nucleus ${}^{39}\mathrm{Mg}$ reports an indication of a low-lying resonance at $0.129$ MeV with $J^{\pi}$ either $7/2^{-}$ or $3/2^{-}$ [59]. However, due to the unavailability of experimental data, this does not rule out the existence of other possibilities such as a single resonant structure or more than two states, and in any case, the spin-parity assignment is not unambiguous.
Table 1: Parameter sets for the ${}^{38}\mathrm{Mg}+n$ (upper-panel) and ${}^{37}\mathrm{Na}+n$ (lower-panel) Woods-Saxon interactions, Eq. (6). Here $a$ is diffuseness, $V^{(l)}_{0}$ is the potential depth and $E_{R}$ is the position of the resonances. Note that $r_{0}=1.25$ fm, and $V_{ls}=16.842$ MeV (for ${}^{38}\mathrm{Mg}+n$) and $16.324$ MeV (for ${}^{37}\mathrm{Na}+n$) are fixed.
@@
$V_{{\rm core}+n}(r)=\left(-V^{(l)}_{0}+V_{ls}\lambda_{\pi}^{2}\bm{l}\cdot\bm{s}\frac{1}{r}\frac{d}{dr}\right)\frac{1}{1+\exp[(r-R_{c})/a]}\,,$ (6)
where $V^{l}_{0}$ is in general $l$-dependent, and $R_{c}=r_{0}A_{c}^{1/3}$ with $A_{c}$ the mass number of the core. The spin-orbit interaction is written in terms of the Compton wavelength $\lambda_{\pi}=1.414$ fm. Following Ref. [60], the spin-orbit strength is taken to follow the systematic trend [61] and has the value $V_{ls}=16.842$ MeV for ${}^{38}\mathrm{Mg}+n$ and $16.324$ MeV for ${}^{37}\mathrm{Na}+n$. The value $r_{0}=1.25$ fm is as originally suggested for ${}^{31}\mathrm{Ne}$  [60, 30], and we examine three scenarios, see Table 1. In Set $1$ and $2$, $V_{0}$ is chosen to be $l$-independent. In set $1$ it is adjusted to fix the $f_{7/2}$ ground-state resonance of ${}^{39}\mathrm{Mg}$ at $0.129$ MeV, corresponding to the prediction of Ref. [59]. This set gives a “normal” shell-model scenario, with an additional $p_{3/2}$ resonance appearing at higher energy, $0.349$ MeV.
@@
These potentials generate Pauli forbidden single-particle states, which would result in unphysical eigenstates of the three-body Hamiltonian, and need to be removed. In this work, this is achieved through a supersymmetric transformation [62, 63].
@@
As we do not have either theoretical or experimental predictions and data for ${}^{38}\mathrm{Na}$, we use the same potential parameters as for ${}^{39}\mathrm{Mg}$. The only changes are $R_{c}$ and the spin-orbit strength, as explained above. The latter takes into account effectively the tensor force between the unpaired proton and valence neutron [64, 65]. Indeed, since ${}^{39}\mathrm{Na}$ has an odd proton number, its ground state angular momentum is not zero. However, we have found before [30, 66, 58] that approximating the ground state of these light systems as a $0^{+}$ state gives reasonable results. The lower half of Table 1 shows the parameter sets for the ${}^{37}\mathrm{Na}+n$ potential. Our predictions for the low-lying spectrum of ${}^{38}\mathrm{Na}$ show no “normal” ordering of the single-particle states; most scenarios present an inversion, indicating a shell-evolution effect, that we describe effectively by just changing the spin-orbit strength.
4 Three-body ($\text{core}+n+n$) calculations